# Towards Trustworthy Multimodal Moderation via Policy-Aligned Reasoning and Hierarchical Labeling

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Despite recent progress, MLLM-based moderation systems still struggle with several fundamental challenges in real-world moderation tasks.  Challenge 1: Misalignment between model outputs and platform policies. Existing models are trained on annotated data, where annotations are based on human interpretations of complex platform guidelines, which are often expressed as intricate textual rules with edge cases and subjective boundaries. Therefore, while the model learns statistical patterns from labels, it lacks direct access to the underlying rules those labels are based on, which often causes the predictions of models to diverge from current moderation expectations. Challenge 2: Lack of transparency in the decision-making process. Most existing approaches produce black-box predictions without surfacing verifiable evidence or logical explanations (Doshi-Velez and Kim, 2017a; Jhaver et al., 2019), limiting trust and hampering operational collaboration between algorithm teams and human reviewers (Doshi-Velez and Kim, 2017b). Unable to determine whether the model is reliable and which parts of the system need to be optimized, departments disagreed, resulting in significant communication and operations costs.  Challenge 3: Difficulty in modeling fine-grained label distinctions under complex, ambiguous, and adversarial conditions. Real-world moderation scenarios often involve subtle differences between sibling categories that share higher-level semantics but differ in definition boundaries. Misclassifications frequently occur between such closely related categories, e.g., Excessive adultification of minors vs. Inappropriate attire for minors. Worse still, these mistakes between adjacent categories can lead to either over-flagging or under-enforcement, reducing the fairness and reliability of moderation outcomes.
@@
To address these challenges, we propose Hi-Guard, a novel hierarchical content safety framework designed for the decision layer. Hi-Guard incorporates three core innovations.  Innovation 1: Hi-Guard learns from rules, not just data. Instead of learning implicitly from labeled data alone, Hi-Guard explicitly integrates category-level policy definitions into the model prompt. This allows the model to reason over rules at inference time and align its decisions with current platform standards, rather than merely mimicking past annotations. As a result, Hi-Guard can generalize better to unseen cases, reflect the intention of updated policies, and provide traceable justifications for each classification. Innovation 2: Hi-Guard introduces an explicit chain-of-thought (CoT) (Wei et al., 2022) reasoning component, coupled with reinforcement learning from verifiable accuracy and format rewards (RLVR) (Guo et al., 2025b; Team et al., 2025), where the model outputs not only its decision but also a detailed explanation (Kojima et al., 2022). This RLVR-driven policy optimization Group Relative Policy Optimization (GRPO) (Shao et al., 2024), enables Hi-Guard to deliver transparent, auditable, and rule-compliant decisions that can be understood and trusted by human moderators.  Innovation 3: Hierarchical taxonomy and path-based soft reward restraint. We introduce a hierarchical taxonomy in which each category is represented as a structured path spanning domain, topic, subtype, and behavior, mimicking real-world moderation policy granularity. Unlike the binary hard reward in RLVR, a soft-margin reward is introduced to penalize sibling-category confusion more than distant errors, with increasing penalties at deeper levels. This structure-aware design improves the model’s ability to distinguish subtle risks and generalize to ambiguous or unseen cases.
These key innovations together establish a novel decision paradigm: the decision & evidence layer. This new layer replaces the original decision layer, enabling models to output not only classification results but also interpretable rationales that align with platform policies, fostering transparency and rule consistency. Extensive offline and online evaluations demonstrate that Hi-Guard, built upon this paradigm, significantly improves classification accuracy, reasoning quality, and system efficiency, effectively breaking the vicious cycle between models, human reviewers, and moderation rules.  This paper makes the following key contributions:
@@ 2. Related Work
Reinforcement Learning.  To align language model behavior with human preference or task-specific constraints, reinforcement learning has emerged as a popular optimization strategy beyond traditional supervised fine-tuning. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) has been widely used to train LLMs, where scalar preference scores are used to optimize the model policy (Ziegler et al., 2019; Glaese et al., 2022; Wu et al., 2021).  Beyond preference comparisons, many approaches transform human feedback, including pairwise judgments, binary correctness, and rule compliance, into reward signals that supervise the training process (Rafailov et al., 2023; Stiennon et al., 2020; Hancock et al., 2019; Zhou and Xu, 2020; Perez et al., 2019; Sun et al., 2023; Yu et al., 2024; Zhao et al., 2023).  More recently, RLVR (Guo et al., 2025a) has been introduced to incorporate structured and interpretable reward signals, such as classification correctness and format consistency (Liu et al., 2025; Yu et al., 2025).  These reward designs often encourage structured outputs via CoT prompting (Wei et al., 2022; Kojima et al., 2022), enabling models to expose their reasoning process and improving decision transparency. To further enhance training robustness, GRPO (Shao et al., 2024) is introduced, which aggregates policy updates across example groups, improves sample efficiency, and reduces variance under sparse reward conditions.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783934isbn: 979-8-4007-2258-5/2026/08ccs: Security and privacy Social network security and privacyccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processingccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
Despite recent progress, MLLM-based moderation systems still struggle with several fundamental challenges in real-world moderation tasks.  Challenge 1: Misalignment between model outputs and platform policies. Existing models are trained on annotated data, where annotations are based on human interpretations of complex platform guidelines, which are often expressed as intricate textual rules with edge cases and subjective boundaries. Therefore, while the model learns statistical patterns from labels, it lacks direct access to the underlying rules those labels are based on, which often causes the predictions of models to diverge from current moderation expectations. Challenge 2: Lack of transparency in the decision-making process. Most existing approaches produce black-box predictions without surfacing verifiable evidence or logical explanations (Jhaver et al., 2019), limiting trust and hampering operational collaboration between algorithm teams and human reviewers (Doshi-Velez and Kim, 2017). Unable to determine whether the model is reliable and which parts of the system need to be optimized, departments disagreed, resulting in significant communication and operations costs.  Challenge 3: Difficulty in modeling fine-grained label distinctions under complex, ambiguous, and adversarial conditions. Real-world moderation scenarios often involve subtle differences between sibling categories that share higher-level semantics but differ in definition boundaries. Misclassifications frequently occur between such closely related categories, e.g., Excessive adultification of minors vs. Inappropriate attire for minors. Worse still, these mistakes between adjacent categories can lead to either over-flagging or under-enforcement, reducing the fairness and reliability of moderation outcomes.
@@
To address these challenges, we propose Hi-Guard, a novel hierarchical content safety framework designed for the decision layer. Hi-Guard incorporates three core innovations.  Innovation 1: Hi-Guard learns from rules, not just data. Instead of learning implicitly from labeled data alone, Hi-Guard explicitly integrates category-level policy definitions into the model prompt. This allows the model to reason over rules at inference time and align its decisions with current platform standards, rather than merely mimicking past annotations. As a result, Hi-Guard can generalize better to unseen cases, reflect the intention of updated policies, and provide traceable justifications for each classification. Innovation 2: Hi-Guard introduces an explicit chain-of-thought (CoT) (Wei et al., 2022) reasoning component, coupled with reinforcement learning from verifiable accuracy and format rewards (RLVR) (Guo et al., 2025; Team et al., 2025), where the model outputs not only its decision but also a detailed explanation (Kojima et al., 2022). This RLVR-driven policy optimization Group Relative Policy Optimization (GRPO) (Shao et al., 2024), enables Hi-Guard to deliver transparent, auditable, and rule-compliant decisions that can be understood and trusted by human moderators.  Innovation 3: Hierarchical taxonomy and path-based soft reward restraint. We introduce a hierarchical taxonomy in which each category is represented as a structured path spanning domain, topic, subtype, and behavior, mimicking real-world moderation policy granularity. Unlike the binary hard reward in RLVR, a soft-margin reward is introduced to penalize sibling-category confusion more than distant errors, with increasing penalties at deeper levels. This structure-aware design improves the model’s ability to distinguish subtle risks and generalize to ambiguous or unseen cases. These key innovations together establish a novel decision paradigm: the decision & evidence layer. This new layer replaces the original decision layer, enabling models to output not only classification results but also interpretable rationales that align with platform policies, fostering transparency and rule consistency. Extensive offline and online evaluations demonstrate that Hi-Guard, built upon this paradigm, significantly improves classification accuracy, reasoning quality, and system efficiency, effectively breaking the vicious cycle between models, human reviewers, and moderation rules.  This paper makes the following key contributions:
@@ 2. Related Work
Reinforcement Learning.  To align language model behavior with human preference or task-specific constraints, reinforcement learning has emerged as a popular optimization strategy beyond traditional supervised fine-tuning. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) has been widely used to train LLMs, where scalar preference scores are used to optimize the model policy (Ziegler et al., 2019; Glaese et al., 2022; Wu et al., 2021).  Beyond preference comparisons, many approaches transform human feedback, including pairwise judgments, binary correctness, and rule compliance, into reward signals that supervise the training process (Rafailov et al., 2023; Stiennon et al., 2020; Hancock et al., 2019; Zhou and Xu, 2020; Perez et al., 2019; Sun et al., 2023; Yu et al., 2024; Zhao et al., 2023).  More recently, RLVR (Guo et al., 2025) has been introduced to incorporate structured and interpretable reward signals, such as classification correctness and format consistency (Liu et al., 2025; Yu et al., 2025).  These reward designs often encourage structured outputs via CoT prompting (Wei et al., 2022; Kojima et al., 2022), enabling models to expose their reasoning process and improving decision transparency. To further enhance training robustness, GRPO (Shao et al., 2024) is introduced, which aggregates policy updates across example groups, improves sample efficiency, and reduces variance under sparse reward conditions.
@@ 5. Conclusion
Acknowledgements.
This work was supported by the National Key Research and Development Program of China under Grant 2024YFF0509700, National Natural Science Foundation of China (62471290, 62331014), and the Fundamental Research Funds for the Central Universities.